OpenStudy (anonymous):

Implicit differentiation: xy^2 + 3x^2 - y^2 + 15 = 0 at point (-1,3)

OpenStudy (anonymous):

Put x, solve for dy/dx

OpenStudy (anonymous):

Ah, I was doing x(2y)(dy/dx)!

OpenStudy (anonymous):

\[2xy \frac{dy}{dx}+y^2 -6x -2y \frac{dy}{dx}=0\]

OpenStudy (anonymous):

Oh wow, nevermind. *faceplam* Those two are the same, I had a brainfart for a second. Sorry about that, thank you for your response.